OpenStudy (anonymous):

Patti’s monthly pay is $3460.96. She usually uses 3/8 of that amount to pay her household expenses, and 1/4 of that amount to pay for tuition. She deposits the rest into her savings account. This month, Patti deposited twice the amount of money into her savings account than she normally deposits. How much did she deposit into her savings account this month?

OpenStudy (anonymous):

$2595.72

OpenStudy (anonymous):

could you explain how you got plz...if you mind @Musics_life

OpenStudy (anonymous):

Yes. 3460.96*(3/8)=1297.86 3460.96*(1/4)=865.24 1297.86+865.24=2163.1 3460.96-2163.1=1297.6 1297.6*2=2595.72

OpenStudy (anonymous):

i totally understand now! Thanks so muck @Musics_life

OpenStudy (anonymous):

much lol

OpenStudy (anonymous):

Lol, You're welcome!